OpenStudy (anonymous):

a=<3,0,-1>, find a vector b such that comp(sub a) b=2.

OpenStudy (anonymous):

\[comp _{a}b=\frac{ a.b }{ |a| }\]

OpenStudy (anonymous):

comp=2, so I made a vector b=<a,b,c> 2=(3a-c)/sqrt(10), now what?

OpenStudy (anonymous):

\[ b=<0,1,-2\sqrt{10}>\\ a.b = 2\sqrt{10}\\ |a|=\sqrt{10}\\ \]

OpenStudy (anonymous):

\[\frac { a. b} {|a|}=\frac { 2 \sqrt{10}} {\sqrt{10}}=2 \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

where did you get the values for b?

OpenStudy (anonymous):

@eliassaab I understand how you get to a.b=2sqrt(10) but now how you get those values for vector b

OpenStudy (anonymous):

I reread the question, i guess there are many other vectors, so you just pick them as long as they get to 2sqrt(10) by a.b

OpenStudy (anonymous):

There infinitely many vectors that satisfy your condition. I guessed it.\[ |a|=\sqrt {10}\\ a. b = 2 \sqrt {10} \] you find b so that the last equation holds.

OpenStudy (anonymous):

You could have taken \[b=<0, 100, -2\sqrt{10}>\]

OpenStudy (anonymous):

got it, thanks

OpenStudy (anonymous):

yw